OpenStudy (anonymous):

A pair of equations is shown below: y = 3x − 5 y = 6x − 8 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. (5 points) Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer. (5 points)

OpenStudy (anonymous):

THat's the same as saying 3x-5 = 6x - 8. Solve for x, giving you x = 1. Now you plug in x, and get y, which is y = -2. In order to graph it, you would just use the two values you have now for x and y, and that would be your point that they intersect. (1, -2)

OpenStudy (anonymous):

y = 3x - 5 y = 6x - 8 --- put the system of linear equations into standard form --- 3x - y = 5 6x - y = 8 --- copy and paste the above standard form linear equations in to this solver: https://sooeet.com/math/system-of-linear-equations-solver.php --- solution by matrix method: x = 1 y = -2 the lines intersect at the point (1,-2) by definition, the solution of a system of linear equations is the point of intersection (if any) of the lines --- solve by elimination: --- 3x - y = 5 (6x - y = 8)*(-1) --- 3x - y = 5 -6x + y = -8 --- add the two equations: --- -3x = -3 x = 1 y = 3x - 5 y = 3(1) - 5 y = -2 ---

OpenStudy (anonymous):

Thank you guys!

OpenStudy (anonymous):

welcome @dfwkem5601

OpenStudy (anonymous):

You're welcome! @dfwkem5601